OpenStudy (anonymous):

The height of a trapezoid is 8 inches, one of the bases measures 24 inches. If the area of the trapezoid is 160 square inches, find the length of the other base.

OpenStudy (anonymous):

The equation is 24 + x times 8 / 2 = 160 ? right?

OpenStudy (raden):

The area of trapezoid = 160 1/2 * (24+x) * 8 = 160

OpenStudy (raden):

yes, solve for x

OpenStudy (anonymous):

I always get confused when I start solving with these kinds of equations..

OpenStudy (raden):

1/2 * (24+x) * 8 = 160 look that 1/2 * 8 = 4, u have 4 (24+x) = 160 divided by 4 to both side, giving us 4/4 * (24+x) = 160/4 24 + x = 40

OpenStudy (raden):

it looks easier now :)

OpenStudy (anonymous):

x = 16?

OpenStudy (raden):

yep